How much would a 190 kg man weigh near the photosphere of the Sun, where the acceleration due to gravity is 274 m/s2?
1
Expert's answer
2020-12-15T11:33:01-0500

The weight of the man can be found from the formula:


"W=mg=190\ kg\cdot 274\ \dfrac{m}{s^2}=52060\ N."

Answer:

"W=52060\ N."